Jeanne and Harold Kimura want to set up a TDA that will generate sufficient interest at maturity to meet their living expenses, which they project to be $1,200 per month. (Round your answers to the nearest cent.) (a) Find the amount needed at maturity to generate $1,200 per month interest if they can get 7 % interest compounded monthly. (b) Find the monthly payment that they would have to put into an ordinary annuity to obtain the future value found in part (a) if their money earns 9 % and the term is twenty-five years. Part a): ISelect Part b): [Select 4 Previous
Answer #1

1.

amount needed at maturity=1200/(7.25%/12)=198620.7

2.

term 25 years or 25*12=300 months

rate 9.75%

monthly annuity

monthly payment needed:

Using financial calculator: I/Y=9.75%/12 FV=198620.7 N=12*25=300 PV=0 CPT PMT=156.19

In Excel=PMT(9.75%/12,12*25,0,198620.7)=156.19